# Towards Human-centered Explainable AI: A Survey of User Studies for Model Explanations

### v4

@@ 6 A Guideline for XAI User Study Design
[…]
@@
From Table V, we can see that previous works have frequently struggled to prove the effectiveness of XAI even with respect to a control group that is without explanation. When only different explanation techniques are considered, there will always be one winner explanation, but the overall benefit will remain undisclosed (see examples in Appendix LABEL:app-sec:appfindings). Therefore, it is important to compare with a baseline without explanations to rigorously show the strength of XAI. When a comparative design is explicitly desired, baselines such as random explanations [28, 62, 41]).
@@ Appendix B Models and Explanations in XAI User Studies
[…]
[anchor=north, align=center, text width=3cm] at (s0.south) LRP [225] ;
[anchor=north, align=center, text width=3cm] at (s1.south) LIME [147]
[anchor=north, align=center, text width=3cm] at (s2.south) SHAP [148]
GradCAM [149]
IG [213]
SmoothGrad [226] ;
[anchor=north, align=center, text width=3cm] at (s3.south) INN [227]
TCAV [160] ;
[anchor=north, align=center, text width=3cm] at (s4.south) ConceptSHAP [61]
ProtoNet [228] ;
[anchor=north, align=center, text width=3cm] at (s5.south) MAME [229]
Dr.XAI [155]
[…]
VIBI [42]
CLUE [51] ;
@@ Appendix C Measurement Details
[…]
! Tasks Tabular Image/Video Text Other forward simulation [40, 53, 49] [51, 55, 35] [39, 12, 22] [13, 56, 52] [54, 58, 59] [43, 60][55, (VQA)] [40, 50] [46] (Audio) marginal feature effects [48, 40, 47] [12, 22] [40] manipulation / counterfactual sim. [22, 12, 45] [32] [50] feature importance [47, 39, 12] [57, 160] failure prediction [58] relative simulation (selection) [35, 22] other [48] (mental model faithfulness) [56] (class-wise acc.)
@@
[…]
! Studied Paper Metric Definition Source Detail Observed [1] Weight of Advice (WOA) - Degree to which the algorithmic suggestion influences the participant’s estimate. [9, 12, 11, 25, 15, 26] Agreement rate - Percentage of cases in which participants agree with the model. [12] defines the appropriate trust, overtrust and undertrust. [11] defines as adherence Self-reported [2, 3, 21] Trust in Automation [159] On the 7-point Likert scale. [2] adapts the questions. [7, 23] General trust in technology [161] On the 5-point Likert scale. [5] Human-Computer Trust [231] On the 7-point Likert scale. [5] adapts the questions. [8] Trust-TAM (Technology Acceptance Model) [232] On the 7-point Likert scale. [8] includes other self-designed questions. [22] Trust in human-machine systems [233] On the 7-point Likert scale. [28] Unified Theory of Acceptance and Use of Technology Model (UTAUT) [234] On the 5-point Likert scale. [10] Human-Robot Collaborative Fluency Assessment [125] On the 7-point Likert scale [23] Trusting beliefs and intentions [235] On the 7-point Likert scale. [6, 13, 27, 11, 30, 29] [16, 17, 18, 19, 24, 26] Self-designed questionnaire - [6, 11, 17] are on the 7-point Likert scale. [13, 27, 18, 19, 24] are on the 5-point Likert scale. [16] rates from 0 to 100. [16, 13, 18, 19, 24, 30, 29] measure one-dimensional trust. [4] Semi-structured interview -
@@ C.2 Usability
[…]
! Studied Paper Metric Definition Source Detail Workload [21, 16, 3, 66, 31] NASA TLX [163] [48] Memory Performance - Helpfulness [13, 56, 65] [67, 46, 48] Self-designed questionnaire - [13, 56, 65] are on 5-point Likert scale [48, 67, 46] are on 7-point Likert scale [68] Rating - Rating from 1 to 5 [45] Comparison - Users select the most helpful method Satisfaction [16, 18, 30, 29, 19, 6, 70] Self-designed questionnaire - [18, 19, 70] are on 5-point Likert scale [29, 30, 6] are on 7-point Likert scale [16] rates from 0 to 100 [7, 69] User experience of recommendation system [161] [7] adapts the questions on the 5-point Likert scale [69] adapts the questions on the 7-point Likert scale [47, 1] Explanation Satisfaction Scale [164] [1] are on 5-point Likert scale [47] are on 6-point Likert scale Undesired behavior detection [71] Number of identified bugs - Questions about bug identification and solutions [24] Accuracy (percentage of correct answers) and time - Task is to determine the correctness of model answers [53] Deviation between human’s and model’s predictions - Model’s predictions are buggy and human’s predictions should be different. [57] Accuracy (percentage of correct answers) - Task is to identify (ir)relevant features [72] Accuracy of answers - Task is to detect model biases or discrimination [27, 38, 2, 74, 78, 79] Rating - [27]: to judge whether they receive enough information to judge the model process is unfair or not; The other judge the model is unfair or not. [73] Rating - Rating on the unfairness of features Ease of use and others [48, 81, 65, 24, 66] Self-designed questionnaire - [65, 24] are on 5-point Likert scale [48, 81] are on 7-point Likert scale [83] AVAM and UEQ-S [236, 237] Autonomous Vehicle Acceptance Model Questionnaire (AVAM) [236] User Experience Questionnaire-Short (UEQ-S) [237] Both on the 7-point Likert scale [32] Single Ease Question (SEQ) [238] On the 7-point Likert scale [71] User Engagement Scale (UES) [239] On the 7-point Likert scale [37, 82] System Causability Scale [165] On the 5-point Likert scale [3] System Usability Scale [240] On the 5-point Likert scale [85, 20] semi-structured interview - -
@@ Appendix D Findings
[…]
.9! Other Evaluation Dimensions Positive Non-positive / Mixed Trust [2]: balanced training data, [17]: high model performance [23]: high quality of explanations [27]: high AI literacy [29]: interactivity [15]: model confidence [2]: user expertise, insignificant [18]: personal characteristics, insignificant [30]: different topic modeling approaches, insignificant [5]: self-referential pronoun “I” in explanations, negative [22]: user technical literacy, insignificant Understanding Obj. [32]: disentanglement of gen. model [22]: interactivity [45]: ExpO regularization of the model [32, 53]: high dimensionality, negative [47]: contextualization, insignificant [13]: inductive vs. deductive explanations, insignificant [50]: different ML models, insignificant [52]: user expertise, insignificant [58]: instant feedback, insignificant [56]: timing of model errors, mixed Sub. [32]: disentanglement of gen. model [22]: interactivity [44]: user expertise [40]: model correctness, insignificant [36]: QuickSort, insignificant [35]: test of understanding, negative Usability [32]: significant difference in self-reported difficulty dependent on the generative model [7, 29]: interactivity [66]: Parallel Embeddings [27]: high AI literacy [73]: fair features are “current charges” and “criminal history” [30]: different topic modeling approaches, insignificant [18]: personal characteristics, insignificant for satisfaction [56]: early encounters of system weaknesses lead to lower explanation usage [53]: clear model is less useful in debugging [73]: unfair features are “quality of school life” and “education & school behavior”, etc. Human-AI Collaboration Performance [90]: low model complexity [92, 13]: showing model prediction [10]: explanations are positive for novices’ performance but negative for experts’ [43, 15] Showing predictions, insignificant [15]: model confidence
@@
[…]
.9! Evaluation Dimension: Explanations Effect comparison among different explanations Trust [9]: example-based explanations are positive in trust building [13]: deductive (rule-based) explanations $>$ inductive (example-based) explanations in decision-making tasks, but contrary in proxy tasks [17]: different explanations positively affect different beliefs of trust [19]: proposed explanation interfaces (different visualizations), SCATTER $>$ RANK and SCATTER $>$ TUNER but insignificant [14] HEX-RL (theirs) > LSTM-attention (for RL agents) Understanding Obj. [46]: Cues and Counterfactuals > Saliency (audio data) [48]: Sparse Lin. > COGAM > GAM [49]: MAME > SP-LIME [51]: CLUE > Sensitivity, Human CLUE, Random (for uncertainty) [52]: Natural images > synthetic (activation prediction) [57]: Counterfactuals (INN) = (proposed) Baseline Expl. > Concepts [55] Anchors > LIME Sub. [33]: local+global explanation > local/global explanation [17]: example-based explanations (normative/comparative) improve the subj. understanding [40] LIME $\geq$ Composite, Prototypes and others [37]: closest and plausible counterfactuals, difference insignificant [33]: local+global explanation > local/global explanation [14] HEX-RL (theirs) > LSTM-attention (for RL agents) [44]: visual > textual explanations Usability [48]: sLM $\leq$ COGAM $<$ GAM, insignificant for self-reported cognitive load [47]: contextualizing/exploration improve user’s satisfaction, but no significant impact when interacting both factors [81]: diff. expl. (e.g. local expl., counterfactuals,…) [21]: GAM vs. SHAP, pos. for cognitive load [16]: diff. interfaces, pos. for cognitive load [46]: counterfactual+cues > saliency, pos. for helpfulness [68]: DEAML $>$ EFM (feature-level expl.) $>$ PAV (“people also viewed” expl.) for usefulness in RS [56]: Salient video segments $>$ Confidence scores, Component combinations shown for helpfulness [13]: deductive (rule-based) has higher cognitive load than inductive (example-based) in proxy tasks, deductive (rule-based) $>$ inductive (example-based) in helpfulness in decision-making task [37]: closest and plausible counterfactuals, difference insignificant [69]: text explanation $>$ visual explanations in user experience (e.g., satisfaction) [19]: proposed explanation interfaces (different visualizations), SCATTER $>$ RANK and TUNER $>$ SCATTER in satisfaction, RANK $>$ SCATTER and TUNER $>$ SCATTER in usefulness, but all insignificant [57]: Counterfactuals (INN) = (proposed) Baseline Expl. > Concepts in bias detection [72]: AReS (theirs) > AR-LIME [74]: sensitivity- and case-based explanations are rated as least fair when they expose a bias of the model [76]: acceptance of the gender-aware career recommender > gender-debiased [75]: significant preference for equalizing false positives over equalizing accuracy [27]: the amount of information positively relates with perceived fairness [2]: data-centric explanations that indicate balanced training data raise the fairness rating Human-AI Collaboration Performance [13]: both deductive (rule-based) explanations and inductive (example-based) explanations are positive, no significant difference

### v5

@@ 6 A Guideline for XAI User Study Design
                                                                                                                                        Before the User Study Measured Quantity measurable quantity according to a general definition or application-based measure conceptualize measures in social sciences and in the context of XAI Control Group include baseline without explanations include baseline with other (e.g., random) explanations in case of a comparative study Proxy Task carefully balanced difficulty close to the anticipated real-world application participants might have some background knowledge of the task but not too much to bias their judgment or performance Measurements measurements are aligned to definitions independent measurements vs. participant-interdependent measurements aggregation of the measured quantities across participants avoid one-dimensional measures pre-register study online to prevent p-hacking run a pre-study and expert interviews for initial feedback on the envisioned measures
                                                                                                                                        During the User Study Experiment Script prepare a script with explicit instructions and procedure of the experiment obtain consent avoid inadvertent cues Participants recruit a representative number of participants for each study subgroup according to the experimental design and required expertise adhere to the real-world distribution of users according to the application at hand Data Quality attention checks during a long survey preventing order and learning effect (within-subject)
                                                                                                                                        After the User Study Statistical Verification comparison between conditions: use ANOVA, t-test, etc. many dependent variables: use SEM, multi-level model, etc. distributional assumption checks for different statistical tests (e.g., normality tests) Compute measures like Cronbach’s $\alpha$ for verifying scales and Fleiß’s $\kappa$ for detecting randomness in the answers of participants or consistent behavior, i.e., agreement beyond chance in their answers Effect Verification correlation between the subjective and behavioral measures agreement or disagreement with expected (e.g., theoretical) results Reporting (aggregated) measurements (across participants) and their statistical validity model and hyperparameters for reproducibility number of participants, characteristics of participants, recruitment exact conditions analysis methods (name and parameters of the test)
@@
From Table V, we can see that previous works have frequently struggled to prove the effectiveness of XAI even with respect to a control group that is without explanation. When only different explanation techniques are considered, there will always be one winner explanation, but the overall benefit will remain undisclosed (see examples in Appendix D). Therefore, it is important to compare with a baseline without explanations to rigorously show the strength of XAI. When a comparative design is explicitly desired, baselines such as random explanations [28, 62, 41]).
@@ Appendix B Models and Explanations in XAI User Studies
2015201520152015201620162016201620172017201720172018201820182018201920192019201920202020202020202021202120212021 LRP [225] LIME [147] SHAP [148] GradCAM [149] IG [213] SmoothGrad [226] INN [227] TCAV [160] ConceptSHAP [61] ProtoNet [228] MAME [229] Dr.XAI [155] SECA [230] VIBI [42] CLUE [51]
@@ Appendix C Measurement Details
Tasks Tabular Image/Video Text Other
forward simulation
[40, 53, 49]
[51, 55, 35]
[39, 12, 22]
[13, 56, 52]
[54, 58, 59]
[43, 60][55, (VQA)]
[40, 50]
[46]
marginal
feature effects
[48, 40, 47]
[12, 22]
[40]
manipulation /
counterfactual sim.
[22, 12, 45] [32] [50]
feature importance [47, 39, 12] [57, 160]
failure prediction [58]
relative simulation
[35, 22]
other
[48]
(mental model
faithfulness)
[56]
(class-wise
acc.)
@@
Studied
Paper
Metric
Definition
Source
Detail
Observed [1] Weight of Advice (WOA) -
Degree to which the algorithmic suggestion
influences the participant’s estimate.
[9, 12, 11, 25, 15, 26] Agreement rate -
Percentage of cases in which participants agree with the model.
[12] defines the appropriate trust, overtrust and undertrust.
[11] defines as adherence
Self-reported [2, 3, 21]
Trust in
Automation
[159]
On the 7-point Likert scale.
[2] adapts the questions.
[7, 23]
General trust in technology
[161] On the 5-point Likert scale.
[5]
Human-Computer
Trust
[231]
On the 7-point Likert scale.
[5] adapts the questions.
[8]
Trust-TAM (Technology
Acceptance Model)
[232]
On the 7-point Likert scale.
[8] includes other self-designed questions.
[22]
Trust in human-machine systems
[233]
On the 7-point Likert scale.
[28]
Unified Theory of Acceptance
and Use of Technology Model (UTAUT)
[234] On the 5-point Likert scale.
[10]
Human-Robot Collaborative
Fluency Assessment
[125] On the 7-point Likert scale
[23]
Trusting beliefs
and intentions
[235] On the 7-point Likert scale.
[6, 13, 27, 11, 30, 29]
[16, 17, 18, 19, 24, 26]
Self-designed
questionnaire
[6, 11, 17] are on the 7-point Likert scale.
[13, 27, 18, 19, 24] are on the 5-point Likert scale.
[16] rates from 0 to 100.
[16, 13, 18, 19, 24, 30, 29] measure one-dimensional trust.
[4]
Semi-structured
interview
@@ C.2 Usability
Studied
Paper
Metric
Definition
Source
Detail
Workload [21, 16, 3, 66, 31] NASA TLX [163]
[48] Memory Performance -
Helpfulness
[13, 56, 65]
[67, 46, 48]
Self-designed
questionnaire
[13, 56, 65] are on 5-point Likert scale
[48, 67, 46] are on 7-point Likert scale
[68] Rating - Rating from 1 to 5
[45] Comparison - Users select the most helpful method
Satisfaction [16, 18, 30, 29, 19, 6, 70]
Self-designed
questionnaire
[18, 19, 70] are on 5-point Likert scale
[29, 30, 6] are on 7-point Likert scale
[16] rates from 0 to 100
[7, 69]
User experience of
recommendation system
[161]
[7] adapts the questions on the 5-point Likert scale
[69] adapts the questions on the 7-point Likert scale
[47, 1]
Explanation
Satisfaction Scale
[164]
[1] are on 5-point Likert scale
[47] are on 6-point Likert scale
Undesired behavior detection [71]
Number of
identified bugs
- Questions about bug identification and solutions
[24]
Accuracy (percentage of
correct answers) and time
Task is to determine the
correctness of model answers
[53]
Deviation between human’s
and model’s predictions
Model’s predictions are buggy and
human’s predictions should be different.
[57]
Accuracy (percentage of
correct answers)
Task is to identify
(ir)relevant features
[72]
Accuracy of
answers
Task is to detect model biases
or discrimination
[27, 38, 2, 74, 78, 79]
Rating
[27]: to judge whether they receive enough information
to judge the model process is unfair or not;
The other judge the model is unfair or not.
[73]
Rating
Rating on the unfairness of features
Ease of use and others [48, 81, 65, 24, 66]
Self-designed
questionnaire
[65, 24] are on 5-point Likert scale
[48, 81] are on 7-point Likert scale
[83]
AVAM and UEQ-S
[236, 237]
Autonomous Vehicle Acceptance Model Questionnaire (AVAM) [236]
User Experience Questionnaire-Short (UEQ-S) [237]
Both on the 7-point Likert scale
[32]
Single Ease
Question (SEQ)
[238] On the 7-point Likert scale
[71]
User Engagement
Scale (UES)
[239] On the 7-point Likert scale
[37, 82]
System Causability
Scale
[165] On the 5-point Likert scale
[3]
System Usability
Scale
[240] On the 5-point Likert scale
[85, 20]
semi-structured interview
@@ Appendix D Findings
Other Evaluation Dimensions
Positive Non-positive / Mixed
Trust
[2]: balanced training data,
[17]: high model performance
[23]: high quality of explanations
[27]: high AI literacy
[29]: interactivity
[15]: model confidence
[2]: user expertise, insignificant
[18]: personal characteristics, insignificant
[30]: different topic modeling approaches, insignificant
[5]: self-referential pronoun “I” in explanations, negative
[22]: user technical literacy, insignificant
Understanding Obj.
[32]: disentanglement of gen. model
[22]: interactivity
[45]: ExpO regularization of the model
[32, 53]: high dimensionality, negative
[47]: contextualization, insignificant
[13]: inductive vs. deductive explanations, insignificant
[50]: different ML models, insignificant
[52]: user expertise, insignificant
[58]: instant feedback, insignificant
[56]: timing of model errors, mixed
Sub.
[32]: disentanglement of gen. model
[22]: interactivity
[44]: user expertise
[40]: model correctness, insignificant
[36]: QuickSort, insignificant
[35]: test of understanding, negative
Usability
[32]: significant difference in self-reported
difficulty dependent on the generative model
[7, 29]: interactivity
[66]: Parallel Embeddings
[27]: high AI literacy
[73]: fair features are “current charges”
and “criminal history”
[30]: different topic modeling approaches, insignificant
[18]: personal characteristics, insignificant for satisfaction
[56]: early encounters of system weaknesses
lead to lower explanation usage
[53]: clear model is less useful in debugging
[73]: unfair features are “quality of school life”
and “education & school behavior”, etc.
Human-AI
Collaboration Performance
[90]: low model complexity
[92, 13]: showing model prediction
[10]: explanations are positive for novices’ performance
but negative for experts’
[43, 15] Showing predictions, insignificant [15]: model confidence
@@
Evaluation Dimension: Explanations
Effect comparison among different explanations
Trust
[9]: example-based explanations are positive in trust building
[13]: deductive (rule-based) explanations $>$ inductive (example-based) explanations
in decision-making tasks, but contrary in proxy tasks
[17]: different explanations positively affect different beliefs of trust
[19]: proposed explanation interfaces (different visualizations),
SCATTER $>$ RANK and SCATTER $>$ TUNER but insignificant
[14] HEX-RL (theirs) > LSTM-attention (for RL agents)
Understanding Obj.
[46]: Cues and Counterfactuals > Saliency (audio data)
[48]: Sparse Lin. > COGAM > GAM
[49]: MAME > SP-LIME
[51]: CLUE > Sensitivity, Human CLUE, Random (for uncertainty)
[52]: Natural images > synthetic (activation prediction)
[57]: Counterfactuals (INN) = (proposed) Baseline Expl. > Concepts
[55] Anchors > LIME
Sub.
[33]: local+global explanation > local/global explanation
[17]: example-based explanations (normative/comparative) improve the subj. understanding
[40] LIME $\geq$ Composite, Prototypes and others
[37]: closest and plausible counterfactuals, difference insignificant
[33]: local+global explanation > local/global explanation
[14] HEX-RL (theirs) > LSTM-attention (for RL agents)
[44]: visual > textual explanations
Usability
[48]: sLM $\leq$ COGAM $<$ GAM, insignificant for self-reported cognitive load
[47]: contextualizing/exploration improve user’s satisfaction,
but no significant impact when interacting both factors
[81]: diff. expl. (e.g. local expl., counterfactuals,…)
[21]: GAM vs. SHAP, pos. for cognitive load
[16]: diff. interfaces, pos. for cognitive load
[46]: counterfactual+cues > saliency, pos. for helpfulness
[68]: DEAML $>$ EFM (feature-level expl.) $>$ PAV (“people also viewed” expl.) for usefulness in RS
[56]: Salient video segments $>$ Confidence scores,
Component combinations shown for helpfulness
[13]: deductive (rule-based) has higher cognitive load than inductive (example-based) in proxy tasks,
deductive (rule-based) $>$ inductive (example-based) in helpfulness in decision-making task
[37]: closest and plausible counterfactuals, difference insignificant
[69]: text explanation $>$ visual explanations in user experience (e.g., satisfaction)
[19]: proposed explanation interfaces (different visualizations),
SCATTER $>$ RANK and TUNER $>$ SCATTER in satisfaction,
RANK $>$ SCATTER and TUNER $>$ SCATTER in usefulness, but all insignificant
[57]: Counterfactuals (INN) = (proposed) Baseline Expl. > Concepts in bias detection
[72]: AReS (theirs) > AR-LIME
[74]: sensitivity- and case-based explanations are rated as least fair when they expose a bias of the model
[76]: acceptance of the gender-aware career recommender > gender-debiased
[75]: significant preference for equalizing false positives over equalizing accuracy
[27]: the amount of information positively relates with perceived fairness
[2]: data-centric explanations that indicate balanced training data raise the fairness rating
Human-AI
Collaboration Performance
[13]: both deductive (rule-based) explanations and inductive (example-based) explanations
are positive, no significant difference